Warning: Can't find topic HCC.TabsJun07P

Back to overview

YOU ARE VIEWING BlogEntryPointx2007x06x14x18x22

Powering Tests Sector 78 - Report 14.06.2007

PowerPoint RD2-Q4_14.06.ppt:

Test Report of Thursday 14.06.2007

  • ElQA
    • RQD.A78 and RQF.A78 circuits released by ElQA. Circuits OK for powering as soon as QPS gives green light.
    • MBA and MBB lines tested OK up to 1 kV with all instrumentation connected. Circuit OK for QPS IST

  • RD2.L8, RQ4.L8 and its correctors
    • Powered to nominal (6000 A) during the whole morning. Ramped down to beam operational current (5980 A) in the begining of the afternoon.
    • Keeping RD2.L8 at operational current, the five Q4 correctors available were ramped to nominal current (72 A) RQ4.L8 is powered acording to the procedure attached (see slide) up to 3.6 kA and 2.95 kA in apertures 1 and 2 respectively. No quench was observed.
    • Pushing for the quench: With D2 at 6kA, aperture 1 of Q4 at 3610 and four correctors at nominal
      • Ramp up from 2950A to 3610 at 1 A/s - OK
      • Ramp down from 3610A to 2950A at 5 A/s -OK
      • Ramp up from 2950A to 3610A at 10 A/s -OK
      • Ramp down from 3610A to 2950 at 10 A/s - OK
      • Ramp up from 2950A to 3610A at 10 A/s - OK
    • Magnet still doesnt quench, so we go on:
      • Ramp up Aperture B1 and B2 from 3610A to 2200A at 10 A/s - OK
      • Ramp up Aperture B2 from 2200A to 3610A at 10 A/s - OK
      • Ramp up Aperture B2 from 2200A to 3610A at 10 A/s - OK

  • RQ5.L8 and its correctors
    • Calibration tests in all the converters

  • 600 A Circuits
    • Very useful data was obtained during the morning on the RQ6.R7 quench detectors
    • Generation and analysis of the QPS data advances in parallel with PCC tests for the available 600A circuits in the DFBAO

  • 60 A type test performed but not finished. Half a day more needed.

Planning for Friday 15.06.2007

  • RQF.A78 and RQD.A78
    • Wait for green light from QPS to go ahead with PIC tests. All heaters and quench loop worked OK, 'Switch Open Failure' functionality to be corrected.
    • PIC1 for both circuits as soon as QPS gives green light.
    • AB/PO will use the characterisation values obtained during ElQA to avoid the connection of a 120A converter for PCC.
    • The DC cables will be connected to the leads once PIC1 is completed.

  • RB.A78
    • QPS IST to start today. They will take around 2 days.

  • 600A Circuits in the DFBAO
    • PCC will continue for the circuits available. QPS will take proffit of them to obtain the needed data for improving the filtering of the detector noise.
    • PIC2 for the circuits in which PCC is completed
    • AB/PO will do some measurements after PIC2 (below 20A)

  • Matching Section ML8
    • 8-hour heat run with all the circuits in the matching section at nominal current
    • Squeezing tests for RQ5.L8 and RQ4.L8 can be carried out in parallel

  • 60 A type test will start after QPS IST for RQF and RQD are completed.

  • Cryo-instrumentation
    • The time needed by QPS to implement the corrections in the 600A quench detector software will be used to test the electronica cards from cryo damaged during the ElQA tests.

Topic attachments
I Attachment History ActionSorted ascending Size Date Who Comment
PowerPointppt RD2-Q4_14.06.ppt r1 manage 37.5 K 2007-06-14 - 19:54 MatteoSolfaroli  
Edit | Attach | Watch | Print version | History: r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r4 - 2007-06-15 - AntonioVergara
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login